### v1

# Establishing validated standards for Home and Work location Detection

Abstract
Smartphone location data have transformed urban mobility research, providing unprecedented insights into how people navigate and interact in cities. However, leveraging location data at scale presents methodological challenges. Accurately identifying individuals’ home and work locations is critical for a range of applications, including commuting analysis, unemployment estimation, and urban accessibility studies. Despite their widespread use, home–work detection methods lack a standardized framework that accounts for differing data quality and that is validated against ground-truth observations. This limits the comparability and reproducibility of results across studies and datasets. In this paper, we present HoWDe, a robust algorithm for identifying home and work locations from mobility data, explicitly designed to handle missing data and varying data quality across individuals. Using two unique ground-truth datasets comprising over $5,100$ individuals from more than $80$ countries, HoWDe achieves home and work detection accuracies of up to $97\%$ and $88\%$, respectively, with consistent performance across countries and demographic groups. We examine how parameter choices shape the trade-off between accuracy and user retention, and demonstrate how these methodological decisions influence downstream applications such as employment estimation and commuting pattern analysis. By supporting in-house pre-processing through a transparent and validated pipeline, HoWDe also facilitates the sharing of privacy-preserving mobility data. Together, our tools and findings establish methodological standards that support more robust, scalable, and reproducible mobility research at both individual and urban scales.
Keywords: GPS data, Home and work detection, Human Mobility, Stop-location sequences, Mobility algorithms, Missing data, Urban analytics
1 Introduction
Location data collected from smartphones has become indispensable for human mobility analyses [1, 2]. These datasets allow studying large samples of individuals globally, providing insights on fundamental aspects of human movements [3], epidemic spreading [4, 5, 6, 7], migration patterns [8, 5], transportation modeling [9], urban planning [10, 11], and the impact of natural disasters and societal shocks on population movements [12, 13, 14, 15, 16]. In light of these advancements, developing robust methods for location data analysis has become a pressing need for scientists and practitioners [17].
A critical component of location data analyses involves the detection of individuals’ visits to their residential [18, 19, 20], and work location [17]. The accurate identification of home and work locations is key for mobility research, as it enables the analysis of factors such as socio-economic background [21, 22, 15], commuting behaviors [23, 16], social contact networks [24, 14], and accessibility to amenities and services [25, 14]. More broadly, it contributes to advancements in computational demography and social science research [26, 17].
The most widely used methods for home and work detection are heuristic approaches that leverage the regularity of human behavior, such as the tendency to spend nights at home and working hours at a workplace [27, 28, 29]. These methods face significant challenges. First, they are sensitive to missing data, which is common in smartphone records [17]. Data quality varies across individuals and over time, as location gaps arise due to factors like inactivity, disabled services, or poor signal, introducing biases in home and work inference [30, 31, 17]. Second, they rely on the choice of parameters, such as time span required to establish a home location or the minimum number of hours per day spent at work.
Despite their importance for mobility analyses, the extent to which heuristic approaches for home/work location detection are reliable remains unclear, and there is no consensus on how to choose one specific algorithm and its parameters’ values [21, 32, 33, 34, 35, 17], in part due to a lack of ground truth data. Validation efforts so far have been limited, focusing on small subsets of individuals, home detection alone, and datasets with low spatial resolution and short time periods [34, 35, 32, 21]. Furthermore, the influence of key parameter choices (such as the length of the period considered or the minimum activity level necessary to define home/work locations) has not been understood in relation to algorithmic performance. The lack of thorough validation has hindered the refinement and harmonization of these methods into standardized best practices [36, 17]. In turn, the absence of such standards can impact the comparability and reproducibility of mobility studies [17].
In this work, we leverage two distinct ground truth datasets to address these key challenges in home/work location detection. Both datasets include sequences of visits to detailed locations with corresponding ground truth: the first dataset (D1) contains self-reported home and work locations from $4,599$ users collected between 2015 and 2019, while the second (D2) is based on independent expert annotations for $500$ users throughout 2020 (see Methods for details).
First, we leverage these data to uncover behavioral patterns in home and work visits. This allows us to empirically assess whether the assumptions underlying home and work detection hold in practice. Additionally, it enables us to quantify the variability in home-work visit behavior across a diverse set of users.
Second, we formalize HoWDe, an algorithm for home and work detection that enhances existing heuristics by integrating key elements from previous work into a comprehensive approach. HoWDe combines the explicit detection of both home and work locations [14, 29], the identification of the workplace based on recurrent visits during working hours [29], the identification of the home as the most frequently visited location at night [35, 28, 34, 21], a sliding window to identify home and job changes over time [14], with tunable window length [14]. Additionally, HoWDe improves upon existing approaches by better accounting for data gaps and their variation across users and over time. We provide an open-source implementation of the algorithm [37].
Finally, we assess HoWDe’s performance against the ground truth data. We provide a detailed analysis of the role of its parameters, and we study how performance changes across age groups, gender, countries, and urbanization levels. We conclude by quantifying how errors in the estimation of home and work locations propagate to key downstream tasks such as estimating employment rates or commuting distances.
2 Results
2.1 Unveiling temporal patterns of home-work visits
We analyze the regularity of home and work visits by measuring how often and when individuals are recorded at their annotated locations. Figure 1a-b shows the probability of being at home or work by hour of the day. We find that home visits peak between midnight and 6 AM, while work visits are most frequent during typical business hours (between 9 AM and 4 PM).
Patterns of visits to the home locations are similar between weekdays and weekends (see Figure 1c-e). For the work location, we observe a drastic reduction in the fraction of individual visiting their work location during weekends (see Figure 1d-f).
Results are consistent across both datasets, with overall higher probabilities of being at home (and lower probabilities of being at work) found in D2 due to the COVID-19 pandemic containment policies in place during most of the period covered by the data (see Sec.SI-3).
Our analysis suggests that work location is best determined using weekday daytime hours, while home location can be identified using night hours, regardless of weekdays or weekends, similar to previous approaches [2, 28, 35].
However, while general trends exist, we also find significant variations across days and users, highlighting pronounced behavioral heterogeneities at the individual level. Figure 1c-f reports disaggregated visit patterns for single “user-day” pairs. Most users’ daily patterns can be grouped into six behavioral profiles based on home and three on workday visits (see Methods for details on how behavioral profiles are obtained).
Figure 1: Behavioral visit profiles to home and work locations a. Fraction of visits per hour of the day to the annotated home locations in dataset D1, for weekdays (black) and weekends (grey) across individuals. The shaded orange area highlights the period with the highest probability of being at home. b. Fraction of visits per hour of the day to the annotated work locations in dataset 1, for weekdays and weekends. The shaded blue area highlights the period with the highest probability of being at work during weekdays. c. User-day profiles by hours at home (orange), other locations (grey), and periods without location data (light-grey) during weekdays. d. Hours at work (blue), other location (grey), or without data (light-grey), per hour of the day across days with data for weekdays. e-f. Show user-day profiles during weekends for hours at home (orange) and work (blue) respectively. The user-day profiles are sorted by cluster, hour of the first home visit, total time spent at home, and hours without data. Clusters are separated by black vertical lines, and the fraction of days per cluster is noted on the x-axis.
With respect to home visits, most user days fall into one of three typical profiles: i) days when users mostly stay at home, ii) when they are rarely/never at home, and iii) when they are away from home during typical business hours (see Figure 1c,e). Similarly, work-behavior profiles are described by three main daily patterns: i) days when users spend business hours at work, ii) when they are rarely/never at work, and iii) when they spend night hours at work (Figure 1d,f). We find that the distribution of individuals across these profiles is generally independent of their socio-demographic characteristics, such as age, gender, and urbanization (see Sec.SI-5).
As the behavioral profiles we identified are based on single user-day activity data, an important question is whether individuals switch between these profiles or, rather, a single profile consistently represents them. We find that users visit a median of $5$ out of $6$ home clusters and $2$ out of $3$ work clusters (for both D1 and D2). We measure the individual inter-profile variability through the normalized entropy, a measure ranging from 00 (when all days of a user fall into a single profile) to $1$ (when all profiles are equally important) [38] (see Methods). On average, users exhibit a normalized entropy of $0.66\pm 0.01$ for home-profile changes between weekdays and $0.48\pm 0.01$ for weekends ($0.62\pm 0.01$ and $0.42\pm 0.01$ for D2 respectively), and $0.51\pm 0.01$ for work-profile changes on weekdays only ($0.40\pm 0.02$ for D2). These values are comparable to the ones obtained when randomly allocating profiles among user days, confirming that users do not have a single dominant behavioral profile (see Table si.1).
All in all, our findings suggest that focusing on the same profiles across users is unlikely to favour one specific socio-demographic group, thus supporting the choice of focusing on weekday daytime hours (9 am to 4 pm) for detecting work locations and nighttime hours (12 am to 6 am) for detecting home locations, regardless of weekdays and weekends. These time windows, combined with minimum data coverage requirements, form the foundation of the heuristics implemented by the HoWDe algorithm.
2.2 HoWDe: a Home and Work Detection algorithm
In this section, we introduce HoWDe, a dynamic home and work detection algorithm that can capture residential moves and job changes over time through a sliding window approach. A key feature of HoWDe is the identification of home and work locations based on the proportion of time spent at a given location, rather than the absolute amount of time as approached in previous literature [29, 28, 14, 15, 33, 34]. This enables HoWDe to accommodate users with varying amounts of missing data over time (see Figure 1c-f), through tunable parameters that set data-quality filters. We share an open-source implementation of the algorithm in PySpark [37].
2.2.1 HoWDe workflow
HoWDe takes as input a user’s pre-processed sequence of stops, where each stop is fully defined by a user-location unique ID and a start and end timestamp.
First, we perform a temporal aggregation of the stop sequences into hourly-bin sequences. In each hourly bin, we retain the location where the individual has spent most of the time. This aggregation limits sequences to a maximum of 24 locations per day, boosting computational efficiency and making HoWDe faster than previous approaches based on sliding windows, with only limited information loss [14, 15].
Second, we exclude days with insufficient hourly bins with data (this step is controlled by the tunable parameter $dhn$, see Table 2). This is done separately for work detection (where we consider business hours) and home detection (where we consider night hours, see Figure 2b). Hence, it is possible that one day is retained only for the home detection or for the work detection.
Third, for each day $t$, we calculate the fraction of hourly-bins spent at each location during nighttime (24 pm to 6 am for home detection), and typical business hours (9 am to 4 pm on weekdays, for work detection). To compute such a fraction, we use as a denominator the number of hourly bins with records (Figure 2.b). Unlike previous approaches that rely on absolute time spent at each location, considering fractions can better accommodate users with varying amounts of missing data over time.
Fourth, for each day $t$, we consider a window of size $\Delta_{T}+1$, covering the range included in $t\pm\Delta_{T}/2$ days (see Figure 2.c). This window acts as a memory component for the algorithm to better detect behavior in the period before and after a specific day. In each window, we average the fraction of hours a location was visited (see Figure 2c).
Thus, for each window, we obtain a collection of potential home and work locations. Among these, we retain only those that meet minimum requirements with respect to the proportion of visits within and across days. These requirements are controlled by parameters $dn_{H,W}$, $hf_{H,W}$, $df_{H,W}$ (see Table 2). If no location meets the requirements, no home (or work) is assigned for that specific window.
Finally, we sort the home and work collections in two different ways. For home, we sort based on the typical fraction of hourly bins the location is visited within a day. For work, we sort based on the typical fraction of days the location is visited within the window. If no work location can be selected based on the fraction of days, the sorting is based on the fraction of hourly bins. The top location from these two lists is then assigned as home (or work). By prioritizing the fraction of days over the fraction of hours for work detection, we are emphasizing work as a location defined by regular visits throughout the workweek (see Sec.SI-6).
In summary, the algorithm allows the control of two aspects. It enables the user to specify the criteria for determining when there is enough data to attempt home and work detection. Also, it enables to control of when a location can be considered as a candidate home or work location, all while accounting for the fact that users can have different levels of missing data.
Figure 2: Schematics of the HoWDe algorithm workflow for home/work location detection. a. Network of visited locations. The diameter represents the number of visits. In orange, we report the stops visited during night hours, and in blue are the ones visited during work hours. b. Sequences of locations visited hourly. The pie charts display the allocation of visits across locations during the night hours (orange) and work hours (blue) on each day. Note that on Saturday, 6/01, the night hours data was discarded since there were not enough hours with available data (exemplifying the filter tuned by the parameter $dhn$). In this minimal example, no location appears during both night and work hours; however, this may happen in real data. c. For each day, $t$ we consider a sliding window of $3$ consecutive days centered on $t$. For work detection, the sliding window excludes weekends. d. Aggregation step for the window centered on Friday, 5th. For each location, we compute the average fraction of time it was visited during night hours (orange) and during work hours (blue). Additionally, we compute the fraction of days the location is visited at least once during work hours. Locations are then sorted in descending order based on these fractions, and the top ones are selected as the estimated Home (L1) and Work (L2) locations on Friday.
2.3 Validation against ground truth data
To assess how well the procedure described above can capture individuals’ home and work locations, we measure i) the accuracy of the algorithm as the fraction of home/work locations that are correctly identified, and ii) the fraction of not-detected users, measured as the number of users where a home/work location was not detected against the number of users for which a home/work location was reported in the ground truth data (see Methods for more details). We run the validation using the ground-truth datasets D1 and D2, described in Methods. In dataset D2, annotators reviewed complete stop sequences; therefore, the temporal dimension (changes of home and work locations) can not be considered. For this reason, for D2, we consider a single window covering the entire data period (see Sec.SI-7 for parameter details).
2.3.1 Algorithm performance across parameter choices
We find that sliding window length $\Delta_{T}$ and minimum home-time fraction $hf_{H}$ have the greatest impact on the accuracy of HoWDe’s home detection, while other parameters have a smaller effect (see Sec.SI-8).
Both the accuracy of the home location detection and the fraction of non-detected homes increase as the minimum home-time fraction threshold $hf_{H}$ increases (Figure 3a,b). In dataset D1, the fraction of non-detected homes increases with the window size, but accuracy shows no substantial improvement beyond $\Delta_{T}=28$ days ($t\pm 2$ weeks). This suggests that a window of $4$ weeks offers a good balance between high accuracy and user retention for home detection.
Figure 3: Home and Work validation across configurations a. Accuracy of the home location detection for datasets D1 (first column) and D2 (second column). Each marker represents different configurations of the minimum fraction of hours an individual must be at home per night. For dataset D1, we explore how the results change for each sliding window increase (x-axis), while for dataset D2, we focus on a single sliding window (comprising the entire data period). b. Fraction of not detected of the home location detection for dataset D1 (first column) and dataset D2 (second column), sharing with the panel a) the legend. c. Accuracy of the work location detection for datasets D1 (first column) and D2 (second column). Each marker represents different configurations of the minimum fraction of hours an individual must be at work within the typical business hours, and the minimum fraction of days an individual should visit the potential work across weekdays. For dataset D1, we explore how the results change for each sliding window increase, while for dataset D2, we keep the same sliding window size (comprising the entire period). d. Fraction of not detected of the work location detection for dataset D1 (first column) and dataset D2 (second column). The legend is shared with panel c.
Table 1 reports the results by focusing on two configuration setups: the minimum configuration, which maximizes accuracy while minimizing the number of non-detected users, and the maximum configuration, which prioritizes maximum accuracy at the cost of higher user non-detection rates. We also report the results obtained by two state-of-the-art home/work detection algorithms: Atlas [28, 2] and TimeGeo [29] (see Methods for more details on these algorithms).
The maximum configuration is reached for $\Delta_{T}=28$ days window and $90\%$ minimum home-time fraction (see Sec.SI-7 for parameter details). With this choice of parameters, the algorithm achieves an accuracy of $97.30\%\pm 0.39\%$ for D1, and $88.47\%\pm 1.13\%$ for D2, but comes at the cost of excluding nearly two out of every five users (see Table 1). The minimum configuration is obtained for parameters $\Delta_{T}=28$ days window and $50\%$ home frequency (see Sec.SI-7 for parameter details). Here, the algorithm achieves $92.94\%\pm 0.49\%$ accuracy for D1, and $72.34\%\pm 1.13\%$ for D2, while maintaining a fraction of non-detected of $4.29\%\pm 0.38\%$ for D1 and $10.44\%\pm 1.02\%$ for D2.
All in all, we find that HoWDe achieves higher accuracy than previous heuristic methods [29, 2], partly because it excludes users with low-quality data (see Table 1).
Shifting to work location detection, we find that the fraction of correctly detected work locations increases with the window size until $\Delta_{T}=42$days ($t\pm 3$ weeks), after which accuracy stabilizes while non-detection rates continue to rise (see Figure 3c,d).
In the work location detection, two key parameters are the minimum fraction of hourly bins ${hf_{W}}$ and days ${df_{W}}$ required for a location to be selected as work. Examining the impact of these two parameters (see Figure 3c-d, different colors), we find that stricter thresholds enhance accuracy in D1 but increase non-detection rates. For a larger value of ${hf_{W}}=60\%$, increasing ${df_{W}}$ enables significant accuracy gains. In contrast, for ${hf_{W}}=40\%$, changes in ${df_{W}}$ have a weaker effect. We find similar results for D2, when changing the minimum fraction of days ${hf_{W}}$ and fraction of hours ${hf_{W}}$.
HoWDe achieves an average work detection accuracy of $83.84\%\pm 0.80\%$ and $88.14\%\pm 0.83\%$ for the minimum and maximum configurations in dataset D1 (see Sec.SI-7 for parameter details), outperforming baseline methods (see Table 1). Similarly, for D2, HoWDe achieves high accuracy levels ($81.86\%\pm 2.13\%-74.43\%\pm 2.13\%$), but, also in this case, systematically lower than D1 (see Table 1). This difference in performances may be partly explained by the fact that dataset D2 was collected during the COVID-19 pandemic, when people did not follow traditional mobility patterns.
Accuracy HoWDe min HoWDe max Atlas TimeGeo
Home D1 $\mathbf{92.94\pm 0.49}$ $\mathbf{97.30\pm 0.39}$ $81.96\pm 0.55$ $79.28\pm 0.78$
D2 $\mathbf{72.34\pm 1.13}$ $\mathbf{88.47\pm 1.13}$ $70.77\pm 1.21$ $70.38\pm 1.12$
Work D1 $\mathbf{83.84\pm 0.80}$ $\mathbf{88.14\pm 0.83}$ $60.32\pm 0.80$ $39.51\pm 0.92$
D2 $\mathbf{76.43\pm 2.13}$ $\mathbf{81.86\pm 2.18}$ $69.71\pm 2.17$ $57.04\pm 2.40$
Fraction not detected HoWDe min HoWDe max Atlas TimeGeo
Home D1 $4.29\pm 0.38$ $38.75\pm 0.93$ $4.60\pm 0.31$ $\mathbf{0.09\pm 0.06}$
D2 $10.44\pm 1.02$ $39.94\pm 1.43$ $\mathbf{0.0\pm 0.0}$ $\mathbf{0.0\pm 0.0}$
Work D1 $20.64\pm 0.77$ $39.71\pm 0.93$ $3.45\pm 0.31$ $\mathbf{1.86\pm 0.23}$
D2 $34.40\pm 1.93$ $58.73\pm 1.87$ $\mathbf{0.0\pm 0.0}$ $0.67\pm 0.41$
Table 1: Evaluation of home and work detection across methods and datasets.
2.4 Robustness and biases across configurations
HoWDe’s accuracy comes at the cost of user retention. The exclusion of users may introduce biases toward specific demographic groups. To assess these biases, we measure accuracy and user retention for different groups based on age (self-reported), gender (self-reported), urbanization levels, and country GDP. Results are obtained for dataset D1, using the minimum and maximum parameter configurations (see Sec.SI-7).
The largest accuracy gaps appear in suburban areas and middle-income countries: home detection drops by $3.39$ points in suburban areas (compared to urban areas) and work detection by $9.13$ in middle-income countries (compared to high-income countries). Despite accuracy deviations, home detection accuracy remains within $90\%$ $99\%$ and work detection within $77\%$ $89\%$ across all lower-accuracy subgroups (see Sec. SI-9), aligning with the observed range for D1 and D2.
Nevertheless, even minor detection errors can propagate, compromising the overall reliability of the results. To address this potential issue, we examine how tuning the detection algorithm parameters impacts results in two downstream applications: (i) employment rate estimates in Europe and (ii) commuting pattern characteristics.
2.4.1 Estimating regional employment rates
We estimate the regional employment rate by measuring the fraction of individuals with a detected work location, stable for a minimum of $30$ days. We then investigate how the employment estimates obtained from a “minimum” and “maximum” configurations differ and how much they deviate from official regional statistics [39]. This analysis covers $90$ regions across six European countries, each with a minimum of $1000$ individuals per region.
Figure 4.a shows that both configurations produce similar geographical patterns in regional employment. However, while both exhibit moderate to high Pearson correlation, the maximum configuration has a stronger correlation with reported employment rates but estimates fewer regions due to stricter data coverage thresholds. It also has a higher overall relative error (Figure 4.b, and Figure SI-12, except in Germany, where the trend reverses. This suggests that while the maximum configuration enhances individual accuracy, the minimum configuration may be more suitable for population estimates.
2.4.2 Home-Work Commuting distances
This case study examines whether commuting distance, measured as the great-circle distance between detected home and work locations, varies with population density and whether parameter configurations can influence results. We estimate urbanization levels and population density around users’ home locations using the 2022 GHS settlement model grid at 1 km resolution [40].
Our analysis reveals consistent findings across configurations: commuting distances are longer in rural areas compared to urban areas, with an absolute difference of $3.61\pm 0.04$ km for the “minimum” and $3.85\pm 0.06$ km for the “maximum” configurations (see Figure 4.c). Additionally, commuting distances decrease with increasing population density (see Figure 4.d). However, the maximum configuration results in shorter commuting distances, with the most densely populated areas showing the largest differences between configurations. This highlights the need for robustness checks to ensure consistency across different configurations.
Figure 4: Applications for home and work detection a. Employment rate by province, as reported (left), measured using the configuration with the minimum user loss (middle), and measured with the configuration with the highest accuracy (right). b. Difference between the minimum (light teal) and maximum configurations (dark teal) by country, as the correlation between the reported and selected configuration (top), the relative error (bottom). c. Difference in commuting distance (km) between urban (dashed) and rural (dotted) by configuration. d. Difference in commuting distance (km) against the median and standard error of the population density bins ($\frac{people}{km^{2}}$).
3 Discussion
In this work, we contributed to the development of accurate and robust methods for home and work detection. Using two datasets (D1, D2) covering $\sim 5,100$ individuals from over $80$ countries, we characterized home/work visit patterns. We found that individuals are most likely at home between midnight and 6 AM ($\sim 75\%$ probability) and at work between 9 AM and 4 PM on workdays ($\sim 50\%$). However, visit patterns and missing data vary widely across users and over time, and even within individual users’ routines.
Building on these insights, we provided an open-source implementation of HoWDe, a home/work detection algorithm that integrates existing approaches while improving robustness against missing data. We validated HoWDe against ground truth datasets, achieving up to $\sim 97\%$ accuracy in D1 and $\sim 88\%$ in D2 for home detection, and up to $\sim 88\%$ accuracy in D1 and $\sim 82\%$ accuracy in D2 for work detection.
Our analysis highlights the trade-off between accuracy and retaining users with low-quality data. We identified two key parameters affecting performance: the window length considered and the minimum visit proportion required for classification. A window of at least $28$ days ensures a reliable home estimation, while work windows longer than $42$ days reduce user retention without significant accuracy gains.
We also revealed how the algorithm’s accuracy varies across population groups, with the largest gaps appearing between urban and suburban areas (for home detection), and between high- and middle-income countries (for work detection). Although these accuracy gaps persist across configurations, two case studies show that prioritizing accuracy over user retention is preferable for individual-level tasks, whereas maximizing retention may be more valuable for population-level analyses.
The insights gained from this study, along with a formalized algorithm, can support research on commuting [15, 16], social contact networks [6], and access to amenities [14], with implications for areas such as computational social science, urban planning, and public health [41, 17]. Improving location data processing will enhance our understanding of both universal mobility laws and micro-level social dynamics [42].
Despite its strengths, HoWDe has key limitations. While it allows customization of detection windows and thresholds, more advanced methods may be required to fully capture mobility patterns that are collected by diverse GPS-equipped devices [43]. Its performance may also vary with lower temporal resolution or geographically limited datasets, particularly in underrepresented low-income regions with sparser mobile data. Future work should assess its applicability in rural and less densely populated areas [44].
Detecting home and work locations raises important ethical concerns, particularly regarding privacy [45, 46], as the geographical coordinates of these locations are highly identifying and can enable re-identification of individuals in mobility datasets. Currently, many data holders apply in-house heuristics to detect home and work, but these methods are often opaque, introducing methodological uncertainty. By providing a validated and open-source detection framework, our approach enables data holders to perform this sensitive step internally using a standardized and transparent pipeline. We hope this will facilitate the sharing of privacy-preserving datasets—for example, through the release of obfuscated or aggregated representations—while ensuring methodological consistency and scientific utility for downstream research.
In summary, by advancing methods for home and work detection, this study contributes to a more robust and scalable foundation for mobility research. By providing an open-source, validated solution, we hope to facilitate more reliable and ethical research, contributing to a deeper understanding of human mobility and its implications for society.
4 Material and Methods
4.1 Dataset 1: DTU stop data and self-reported annotations
The first dataset re-uses data collected during 2017-19 already used in previous publications [3], where stop locations were detected using the scalable stop-location detection algorithm InfoStop [47]. The stop location data captures the mobility patterns of 96,065 users between 25 and 64 years old for 81 countries across continents.
The dataset also contains self-recorded home and work labels for $2,309$ labelled home and $2,479$ work locations, collected across $4,599$ mobile phone and smart band users. Note that the labelled dataset is biased towards urban areas and higher- to middle-income countries. The spatial coordinates of locations are no longer available to preserve individuals’ privacy.
All data analysis was carried out in accordance with the European Union’s General Data Protection Regulation 2016/679 (GDPR).
4.2 Dataset 2: Veraset Movements data and annotation
Anonymous raw GPS points are provided at a global level by the Veraset data company in the framework of the World Bank “Monitoring COVID-19 policy response through human mobility data” project. Veraset Movements data consists of raw GPS point information from devices all around the world. In contrast to the majority of GPS data providers, Verasets Movements are sourced from thousands of Software Development Kits (SDK) to reduce sample biases. It features coverage of about $5\%$ of the global population [48].
In our work, unique-device GPS points are processed and analyzed as a proxy of individuals’ mobility and patterns of visits. The data at our disposal spans over a period of $12$ months, starting from Jan 01, 2020, until Dec 31, 2020. In this work, multiple countries were analyzed to assess the impact of the pandemic and the policy responses across a vast spectrum of potentially different behaviors.
Dataset D2, in particular, consists of the GPS trajectories of $100$ distinct devices for each of the $5$ selected countries: Brazil, Colombia, Indonesia, Mexico, and the Philippines. For each country independently, we require $33\%$ of the users to be living in high-wealth neighbourhoods, $33\%$ of the users to be living in medium-wealth neighbourhoods, and the remaining part in low-wealth neighbourhoods. Wealth classification and users’ place of residence are drawn leveraging previous work on GPS mobility in middle-income countries [15]. In particular, each user is also required to have a minimum level of recorded activity, consisting of at least a stop record in $20\%$ in the days before the COVID-19 pandemic declaration and $20\%$ of days with at least a stop record over the entire data period.
This second dataset involves, for each user’s stop location sequences, the expert annotation through detailed inspection of location visits, satellite imagery, and local amenities, focusing on five middle-income countries (Brazil, Colombia, Indonesia, Mexico, and the Philippines) for 500 individuals, from which 287 have at least one work location annotated (see Sec. SI-1 detailed description of the annotation procedure). Given that annotators only labelled the complete longitudinal stop sequences, an explicit consideration of changes in home and work for each individual is not possible.
4.3 Home and Work locations detection algorithms: a primer
Smartphones use an in-built GPS sensor to estimate the device’s latitudes and longitude coordinates over time, at a frequency that varies depending on device usage. The device position is then collected by various location-based apps, such as navigation or social media apps. These data may then be used by location intelligence companies or researchers to study human mobility and its regularities, both at the individual and at the population level. Raw location data utilized for mobility analysis consists of timestamped sequences of (latitude, longitude) records associated with a unique user ID.
The first pre-processing step involved in the analysis is the so-called “stop-location detection” [47, 49]. Stop location detection consists of (i) identifying records that are captured when individuals stay within a small area for a sufficient amount of time (e.g. when they are at home, at the office, or in a restaurant); (ii) applying spatial clustering to all these “static records”, to identify a set of unique locations, i.e. those places where people spend time. After detecting the individuals’ stops-locations, individual trajectories can be summarized as sequences of stays. Each stay is associated with a start time (when the stay started), a duration (how long the stay lasted), and a location ID (where the stay happened).
A second critical processing step often involved in mobility studies is to identify key locations, such as individuals’ home and work locations, starting from the sequences of stays. To achieve this, researchers and practitioners often define heuristics based on the assumption that visits to home and work locations follow specific daily and weekly patterns. While a standard practice does not exist, in this work, we focus on two distinct and widely used state-of-the-art methods that exploit travel circadian rhythms and pattern regularities. These methods were designed and tested over the past decade with different approaches, in a collective scientific effort, separately leveraging diverse sets of mobile-phone-based data (e.g. Call Detail Records (CDRs), eXtended Detail Records, (XDRs), or GPS records) [21, 50, 51, 35] as well as activity-based travel surveys [36]. The main rationale is to rank locations based on the day of the week (weekdays vs. weekends) and the time of day (daytime vs. nighttime) when visits occur.
4.3.1 The Atlas approach
The method used in [2, 6, 52, 28] (hereinafter “Atlas” method) defines a home location as the most probable location visited on weekdays between 10 p.m. and 6 a.m., i.e. the stop location where a device was found spending most of the weekday nighttime. While the Atlas methods mostly focuses on home locations, its definition can be easily generalized to the case of work locations. Following a similar rationale, we define an Atlas work location as the most probable stop location visited during weekdays daytime.
4.3.2 The Time-Geo approach
In contrast with the definition introduced by the Atlas method, more complex heuristics are adopted in Jiang et al. [29] (hereinafter referred to as the “Time-geo” method), in particular for the case of identifying work locations. Time-geo ranks stop locations based on the frequency of visits during weekday nights (between 7 p.m. and 8 a.m.) and weekends and picks the most visited location as an individual’s home. The same logic is applied when detecting the work location, but now focusing on weekday work time boundaries (between 8 a.m. and 7 p.m.) and including additional requirements: i) the candidate’s work location is visited at least three times across the study period, and ii) the distance between the work candidate and an individual’s home location is greater than $500$ meters.
Note that both methods rely on the definition of stop location. Stop location definition might vary depending on the type of data analyzed and the specific characteristics of the data points, e.g. whether you use Call Detail Records (CDR), eXtended Detail Records (XDR), or GPS points. However, the principled choices these algorithms are founded on can naturally be extended to any of these different cases, making them easily generalizable.
In this work, stop locations are defined as spatiotemporally close sets of GPS points. All comparisons between home-work detection methods are always drawn from a set of stop locations constructed following identical procedures. For more information about stop location processing, we refer to the Supplementary Information (SI) Sec.SI-2 and the existing literature [47, 14].
4.4 Home and work visits behavioral clusters
Section 2.1 focused on the temporal attributes of users’ daily mobility patterns. To better highlight groups of users with similar behavioral features, we reported the results in Figure  1.c-f, sorting daily patterns based on the visit hour of the first home visit. behavioral groups were generated using the K-Modes clustering algorithm, a categorical adaptation of the traditional K-Means algorithm [53]. Here, we use daily sequences of 24 hourly bins with three possible stop categories. For example, in the case of home patterns, each sequence consists of either a home (category 1), a visit to “a different location”, e.g., work or recreational locations (category 2), or an hour with missing data (category 3). A similar approach is used for work patterns, where visits to work location are considered as “category 1”, while visits to home location are considered as “a different location” (category 2). The K-Modes algorithm requires as a parameter, the number of desired clusters to compute. In Figure 1.c,e we reported 6 clusters. In Figure 1.d,f we reported 3 and 4 clusters, respectively. Note that, while the number of clusters is the result of the selection of cluster number following the elbow method for each of the four plots, our purpose is to simply highlight the main different visit patterns. Following this idea, in Sec.SI-3, where we present results for the smaller D2 dataset, we manually selected three as the number of clusters used to report the hourly day-visit patterns.
4.5 Behavioral profile entropy
We quantify the degree to which users transition between behavioral profiles measuring the normalized Shannon entropy [38], treating each behavioral cluster as a distinct state, following:
$H_{\text{norm}}=-\frac{1}{\log{k}}\sum_{i=1}^{k}p_{i}\log{p_{i}},$ (1)
where $k$ is the total number of possible clusters, and $p_{i}$ is the probability of occurrence of each state, calculated as: $p_{i}=\frac{n_{i}}{N}$ where $n_{i}$ are the total days classified in cluster $i$ and $N$ the total number of days for a given user. Normalizing by the maximum entropy $\frac{1}{\log{k}}$, ensures that entropy remains comparable across cases (weekend vs. weekday and home vs. work), staying within the same range from 0 (when there is only one profile) to 1 (when there are multiple profiles, and all are equally important).
4.6 HoWDe parameters details
$\mathbf{dhn}$: First, to account for a minimum level of information within a day, we focus only on days with hour sequences consisting of at least three stops, i.e. stops visited in at least three hours of the day (i.e. days with at least three hours of data). The HoWDe algorithm, however, can be tuned to handle different thresholds based on the specific needs or constraints in place.
$\mathbf{dn_{H,W}}$: Second, to account for the potentially varying data quality across different datasets, we introduce two independently configurable pairs of parameters for home and work locations. The first pair of parameters works as a maximum fraction of days without data per sliding window, marking a day with an “undetected home/work location” if the fraction of days without data within the sliding window exceeds the threshold.
$\mathbf{hf_{H,W}}$: A third pair of parameters threshold the minimum fraction of hours per night or day period for a location to be home or work respectively. If the average fraction of visits across all days in the sliding window is below the threshold, the location cannot be considered a potential home/work.
$\mathbf{df_{H,W}}$: Fourth, while home locations are expected to be visited almost every day (provided that enough data are available), work locations are likely to follow a different pattern, as those locations will be subject to local work habits, e.g. four days working week. To account for regularities across days, we introduce an additional requirement for work locations: a work location should be visited for a minimum fraction of days with data.
Table 2: Summary description of the HoWDe configurable parameters. The table reports the parameter names, whether they are used in home detection, work detection, or both, and the description of the parameter purpose. When looking at the hourly bins with data (hours with data), that is the case of $dhn$, $hf_{H}$, and $hf_{W}$, counts and fractions are computed with respect to night/business hourly bins only. Note that, $dhn$, $dn_{H}$, and $dn_{W}$ act as minimum activity thresholds for a given day to have an estimated home or work location estimated. $hf_{H}$, $hf_{W}$, and $df_{W}$ serve a similar purpose but for location-specific visit requirements allowing each location above the threshold to be a candidate home/work location. Specifically, fractions of visits to home/work location candidates are computed relative to the total hours in which an individual has recorded data during the specified period.
Name Usage Description
$\mathbf{\Delta_{T}}$ Home/Work data quality Size of the memory window
$\mathbf{dhn}$ Home/Work data quality Min. number of night/business hourly-bins with data in a day
$\mathbf{dn_{H}}$ Home data quality Max. fraction of days without data in a window
$\mathbf{dn_{W}}$ Work data quality Max. fraction of days without data in a window
$\mathbf{hf_{H}}$ Home estimation Min. fraction of night hourly-bins (avg. over days in the window)
$\mathbf{hf_{W}}$ Work estimation Min. fraction of business hourly-bins (avg. over days in the window)
$\mathbf{df_{W}}$ Work estimation Min. fraction of days within the window
For details on parameter implementation, we refer to the Sec. SI-6, where we provide the algorithm’s pseudocode.
4.7 Accuracy and “fraction of not detected” definition
In this work, we report the performance of the algorithm in terms of two main quantities: “accuracy” and “fraction of not detected”. Here, we provide the definitions of these two concepts. The definitions below are explained in the case of the home detection, and they apply in the same way for the work detection.
The fraction of ‘not detected’ indicates cases where a home location was labelled, but the algorithm couldn’t assign a home due to the thresholds on data quality or minimum fraction of visits. In the case of D1, the dataset consists of a set of user-weeks, corresponding to users that self-reported living in a specific home (the “ground truth home”) in a given week. For some of these user-weeks, our algorithm identifies a “detected home”. We evaluate the performance of the algorithm by comparing pairs of “ground truth home” and “detected home” across user-weeks. The “fraction of not detected” is thus computed as $f_{ND}=1-n_{\bar{H}}/n_{H}$, where $n_{\bar{H}}$ is the number of user-week pairs with a detected home, and $n_{H}$ is the number of user-week pairs with a ground truth home. In the case of D2, where ground truth information is not referring to a specific week, $n_{\bar{H}}$ corresponds to the number of users with a detected home, and $n_{H}$ corresponds to the number of users with a ground truth home.
The “accuracy” captures the fraction of correctly labelled home locations among the set of detected home locations. Therefore, we define accuracy as $Acc=m_{\bar{H}}/n_{\bar{H}}$, where $m_{\bar{H}}$ is the number of user-weeks (for D1) or users (for D2) for which the detected home location matches the ground truth home location. In dataset D2, if annotators identified multiple homes for an individual, we considered it a match if the individual’s detected home location corresponded to at least one of the annotated ground truth home locations. Note that our definition does not correspond exactly to the traditional definition of accuracy because we separately assess the algorithm performance in terms of the detection/non-detection of a home (captured by $f_{ND}$) and the detection of the correct home (captured by $Acc$).
The definitions above are also used in the case of work detection, where only users with a reported work location, $n_{W}$, are included in the denominator of $f_{ND}$, and only those with a detected and a matching work location are counted in the nominator and in the accuracy computation, as $n_{\bar{W}}$ and $m_{\bar{W}}$, respectively.
5 Data and Code Availability
Data to ensure replicability of the analysis performed in this work can be found at
https://doi.org/10.11583/DTU.28846325 [Note: The link will become available upon publication]. To protect user anonymity, we applied the following privacy-preserving transformations to the stop sequences. First, time was discretized into 10-minute intervals. Second, each user’s timeline was shifted so that their first recorded activity occurred in the first week of 1970. Third, entire days were randomly shuffled within each week, while preserving the distinction between weekdays and weekends. All code needed to implement HoWDe is available at [37]. We also provide HoWDe as a Python (PyPI) package at [54].
Acknowledgments
L.L. has been supported by the ERC project “IMMUNE” (Grant agreement ID: 101003183). L.L. acknowledges the support from the “Fondazione Romeo ed Enrica Invernizzi” for the research activities of the ‘Covid Crisis Lab’ at Bocconi University. L.A. and S.D.S. acknowledge support from the Independent Research Fund Denmark (DFF) through the DFF-Research Project 1 (Inge Lehmann) grant “Gender Gaps in Human Mobility”. The authors thank Y. Liao for providing thoughtful feedback on the code release, which helped improve its clarity and usability.
Declaration of Interest statement
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this article.
Author Contributions
S.D.S., L.L., S.F., and L.A. conceived the study. S.D.S., L.L., and O.D.L.C. curated and preprocessed the mobility and ground truth data. Two alternative algorithms were independently developed by L.L., S.F., and O.D.L.C., and by S.D.S. and L.A. The two designs were compared and discussed collaboratively, and the final HoWDe workflow is based on the design by S.D.S. and L.A., with L.L. contributing to its finalization. S.D.S. and L.L. designed the analysis pipeline and conducted the statistical analyses. S.F. and L.A. provided domain expertise and guidance on the interpretation of results. S.D.S. and L.L. drafted the manuscript. All authors discussed the results and approved the final version of the manuscript.